1. The average depth of the gulf of Mexico is three times the depth of the red Sea. the difference between the average depths is 1078 m. what is the average depth of the gulf of mexico and the av erage depth of the red sea?
2. Two of the largest Internet retailers are ebay and amazon.com. Recently, the estimated U.S. sales of ebay were $0.1 billion less than twice the sales of Amazon.com. Given the total sales of $5.6 billion, determin the sales of the eBay and Amazon.com.
3. In a recent year, the best selling DVD was the Lord of the Rings: The Fellowship of the Rings. Another big seller was Harry Potter and the Sorcerer's Stone, which sold 90 million less than the Lord of the Rings. If the two movies totaled 424 million, how many DVD's of each were sold?

Answer by elima(1433)

3. In a recent year, the best selling DVD was the Lord of the Rings: The Fellowship of the Rings. Another big seller was Harry Potter and the Sorcerer's Stone, which sold 90 million less than the Lord of the Rings. If the two movies totaled 424 million, how many DVD's of each were sold?
1)Mexico = M
Red Sea = R
M = 3R
M -R = 1078
replace 3R for M in second equation;
3R - R = 1078
2R = 1078
R=539
Now plug what you got for in the second equation;
M - 539 = 1078
M = 1078 + 539
M = 1617
2)Ebay = E
Amazon = A
E = 2A-.1
E + A = 5.6
Replace 2A - .1 for E in second equation;
2A - .1 + A = 5.6
3A = 5.6 + .1
3A = 5.7
A = 1.9
Now plug that in for A in second equation;
E + 1.9 = 5.6
E = 5.6 - 1.9
E = 3.7
3)L = Lord of the rings
S = Sorcerer's Stone
L + S = 424
S = L - 90
replace L - 90 for S in first equation;
L + L -90 = 424
2L = 424 + 90
2L = 514
L = 257
Now plug that in the first equation:
257 + s = 424
S = 424 - 257
S = 167
:)